# Retrieval, Reasoning, Re-ranking: A Context-Enriched Framework for Knowledge Graph Completion

### v1

@@ Abstract
The Knowledge Graph Completion (KGC) task aims to infer the missing entity from an incomplete triple. Existing embedding-based methods rely solely on triples in the KG, which is vulnerable to specious relation patterns and long-tail entities. On the other hand, text-based methods struggle with the semantic gap between KG triples and natural language. Apart from triples, entity contexts (e.g., labels, descriptions, aliases) also play a significant role in augmenting KGs. To address these limitations, we propose $\text{KGR}^{3}$, a context-enriched framework for KGC. $\text{KGR}^{3}$ is composed of three modules. Firstly, the Retrieval module gathers supporting triples from the KG, collects plausible candidate answers from a base embedding model, and retrieves context for each related entity. Then, the Reasoning module employs a large language model to generate potential answers for each query triple. Finally, the Re-ranking module combines candidate answers from the two modules mentioned above, and fine-tunes an LLM to provide the best answer. Extensive experiments on widely used datasets demonstrate that $\text{KGR}^{3}$ consistently improves various KGC methods. Specifically, the best variant of $\text{KGR}^{3}$ achieves absolute Hits@$1$ improvements of $12.3\%$ and $5.6\%$ on the FB15k237 and WN18RR datasets. The code is anonymously available at https://anonymous.4open.science/r/RRR-KGC-CBA3/.
@@
Muzhi Li2,$1$Equal contribution., Cehao Yang3,$1$Equal contribution., Chengjin Xu1111Equal contribution., Xuhui Jiang1, Yiyan Qi1, Jian Guo1222Corresponding authors., Ho-fung Leung, Irwin King2222Corresponding authors. 1IDEA Research, International Digital Economy Academy 2Department of Computer Science & Engineering, The Chinese University of Hong Kong 3Artificial Intelligence Thrust, Hong Kong University of Science and Technology (Guangzhou) {limuzhi,yangcehao,xuchengjin,jiangxuhui,qiyiyan,guojian}@idea.edu.cn {mzli,king}@cse.cuhk.edu.hk, ho-fung.leung@outlook.com cyang289@connect.hkust-gz.edu.cn
@@ 1 Introduction
Knowledge Graphs (KGs) are graph-structured knowledge bases (KBs) that organize factual knowledge as triples in the form of (head entity, relation, tail entity). Recently, KGs have become a crucial foundation for various downstream applications, such as recommendation systems Wang et al. (2019), question answering Sun et al. (2024), and sentiment analysis Wang and Shu (2023). Nevertheless, mainstream KGs such as Freebase Bollacker et al. (2008) and Wordnet Miller (1995) suffer from serious incomplete issues. This problem highlights the importance of the Knowledge Graph Completion (KGC) task, which aims to predict the missing entity from an incomplete triple.
@@ 2 Related Works
Embedding-based methods are fundamental in KGC, which focuses on learning a set of low-dimensional embeddings for entities and relations with certain geometric or mathematical constraints. Most typically, TransE Bordes et al. (2013) assumes the translated head embedding of a triple are close to the embedding of the tail. DistMult Yang et al. (2015) aims to maximize the Hadamard product of the head, relation, and tail embeddings of each triple. To model symmetric and anti-symmetric relations, ComplEx Trouillon et al. (2016) generalizes DistMult by introducing complex embeddings with Hermitian dot product. RotatE Sun et al. (2019) interprets relation as a rotation operation in complex space, which can effectively infer inversion and composition patterns. ATTH Chami et al. (2020) and GIE Cao et al. (2022) further leverage hyperbolic embeddings and operations to capture the intrinsic hierarchical structure in KGs. Recently, Graph Neural Networks (GNNs) have emerged as powerful methods for graph embedding Zhou et al. (2020). Considering the heterogeneity of KGs, RGCN Schlichtkrull et al. (2018), WGCN Zhao et al. (2021), and CompGCN Vashishth et al. (2020) adapted to consider relation types in their message-passing functions. KBGAT Nathani et al. (2019) proposes a two-layer attentional network to encode each triple and to measure its importance to the tail entity. In addition, NBF-Net Zhu et al. (2021), RED-GNN Zhang and Yao (2022) and A*Net Zhu et al. (2023b) integrate structural information from paths between the two entities, which also supports inductive KGC. Despite the simplicity and high scalability, embedding methods suffer from the long-tail entity distribution in KGs, and hence, cannot generate semantic expressive embeddings for boundary entities, which limits the performance of KGC.
@@ 4 Methods
1) Reasoning (left), 2) Reasoning (middle), and 3) Re-ranking (right).
@@ 5.2 Baselines and Evaluation Metrics
$\textbf{KGR}^{\mathbf{3}}$ (Llama3-8B) 0.536 0.477 0.565 0.636 0.717 0.656 0.759 0.809
@@ 5.3 Main Results
Compared to the selected base KGC models, the corresponding variants of $\text{KGR}^{3}$ consistently boost the performance on both datasets to a large margin. It shows that $\text{KGR}^{3}$ is compatible with various types of KGC models, confirming its strong generality and plug-and-play capability. In contrast, KICGPT fails to outperform the state-of-the-art GNN-based or text-based methods, which underutilizes the power of the LLM. In addition, variants of $\text{KGR}^{3}$ consistently outperform counterparts of the LLM-based baseline DIFT Liu et al. (2024) with the same base KGC models. The performance improvements can be primarily attributed to the incorporation of the reasoning module and the inclusion of entity descriptions for candidate answers. It should also be noted that DIFT necessitates an expensive re-training process and pre-trained KG embeddings to adapt to various KGC models. Additionally, it benefits from SFT data derived from the validation set Liu et al. (2024). Without these additional triples, DIFT cannot guarantee the experimental results as claimed.
@@ 5.4 Ablation Studies
RQ5: How does the volume of SFT data affect the performance of $\text{KGR}^{\mathbf{3}}$?
Figure 3: Hits@$1$ performance over the volume of SFT data on the two datasets with different LLMs.

### v2

@@ Abstract
The Knowledge Graph Completion (KGC) task aims to infer the missing entity from an incomplete triple. Existing embedding-based methods rely solely on triples in the KG, which is vulnerable to specious relation patterns and long-tail entities. On the other hand, text-based methods struggle with the semantic gap between KG triples and natural language. Apart from triples, entity contexts (e.g., labels, descriptions, aliases) also play a significant role in augmenting KGs. To address these limitations, we propose $\text{KGR}^{3}$, a context-enriched framework for KGC. $\text{KGR}^{3}$ is composed of three modules. Firstly, the Retrieval module gathers supporting triples from the KG, collects plausible candidate answers from a base embedding model, and retrieves context for each related entity. Then, the Reasoning module employs a large language model to generate potential answers for each query triple. Finally, the Re-ranking module combines candidate answers from the two modules mentioned above, and fine-tunes an LLM to provide the best answer. Extensive experiments on widely used datasets demonstrate that $\text{KGR}^{3}$ consistently improves various KGC methods. Specifically, the best variant of $\text{KGR}^{3}$ achieves absolute Hits@$1$ improvements of $12.3\%$ and $5.6\%$ on the FB15k237 and WN18RR datasets.
@@
Muzhi Li2,$1$Equal contribution., Cehao Yang3,$1$Equal contribution., Chengjin Xu1111Equal contribution. [^2] Corresponding authors., Xuhui Jiang1, Yiyan Qi1, Jian Guo1222Corresponding authors., Ho-fung Leung [^3] Independent researcher., Irwin King2222Corresponding authors. 1IDEA Research, International Digital Economy Academy 2Department of Computer Science & Engineering, The Chinese University of Hong Kong 3Artificial Intelligence Thrust, Hong Kong University of Science and Technology (Guangzhou) {limuzhi,yangcehao,xuchengjin,jiangxuhui,qiyiyan,guojian}@idea.edu.cn {mzli,king}@cse.cuhk.edu.hk, ho-fung.leung@outlook.com cyang289@connect.hkust-gz.edu.cn
@@ 1 Introduction
Knowledge Graphs (KGs) are graph-structured knowledge bases (KBs) that organize factual knowledge as triples in the form of (head entity, relation, tail entity). Recently, KGs have become a crucial foundation for various downstream applications, such as recommendation systems Chen et al. (2022a, b), question answering Sun et al. (2024), and sentiment analysis Wang and Shu (2023). Nevertheless, mainstream KGs such as Freebase Bollacker et al. (2008) and Wordnet Miller (1995) suffer from serious incomplete issues. This problem highlights the importance of the Knowledge Graph Completion (KGC) task, which aims to predict the missing entity from an incomplete triple.
@@ 2 Related Works
Embedding-based methods are fundamental in KGC, which focuses on learning a set of low-dimensional embeddings for entities and relations with certain geometric or mathematical constraints. Most typically, TransE Bordes et al. (2013) assumes the translated head embedding of a triple is close to the embedding of the tail. DistMult Yang et al. (2015) aims to maximize the Hadamard product of the head, relation, and tail embeddings of each triple. To model symmetric and anti-symmetric relations, ComplEx Trouillon et al. (2016) generalizes DistMult by introducing complex embeddings with Hermitian dot product. RotatE Sun et al. (2019) interprets relation as a rotation operation in complex space, which can effectively infer inversion and composition patterns. ATTH Chami et al. (2020) and GIE Cao et al. (2022) further leverage hyperbolic embeddings and operations to capture the intrinsic hierarchical structure in KGs. Recently, Graph Neural Networks (GNNs) have emerged as powerful methods for graph embedding Zhou et al. (2020); Song et al. (2023b, a), with various applications Ma et al. (2023); Song et al. (2022, 2024) . Considering the heterogeneity of KGs, RGCN Schlichtkrull et al. (2018), WGCN Zhao et al. (2021), and CompGCN Vashishth et al. (2020) adapted to consider relation types in their message-passing functions. KBGAT Nathani et al. (2019) proposes a two-layer attentional network to encode each triple and to measure its importance to the tail entity. In addition, NBF-Net Zhu et al. (2021), RED-GNN Zhang and Yao (2022) and A*Net Zhu et al. (2023b) integrate structural information from paths between the two entities, which also supports inductive KGC. Despite the simplicity and high scalability, embedding methods suffer from the long-tail entity distribution in KGs, and hence, cannot generate semantic expressive embeddings for boundary entities, which limits the performance of KGC.
@@ 4 Methods
1) Reasoning (left), 2) Reasoning (middle), and 3) Re-ranking (right). (Please see Appendix A for detailed prompts.)
@@ 5.2 Baselines and Evaluation Metrics
$\textbf{KGR}^{\mathbf{3}}$ (Llama3-8B) 0.535 0.475 0.564 0.635 0.717 0.656 0.759 0.809
@@ 5.3 Main Results
Compared to the selected base KGC models, the corresponding variants of $\text{KGR}^{3}$ consistently boost the performance on both datasets to a large margin. It shows that $\text{KGR}^{3}$ is compatible with various types of KGC models, confirming its strong generality and plug-and-play capability. In contrast, KICGPT fails to outperform the state-of-the-art GNN-based or text-based methods, which underutilizes the power of the LLM. In addition, variants of $\text{KGR}^{3}$ consistently outperform counterparts of the LLM-based baseline DIFT Liu et al. (2024) with the same base KGC models. The performance improvements can be primarily attributed to the incorporation of the reasoning module and the inclusion of entity descriptions for candidate answers. It should also be noted that DIFT necessitates an expensive re-training process and pre-trained KG embeddings to adapt to various KGC models Liu et al. (2024). Without such a process, DIFT cannot guarantee the experimental results as claimed.
@@ 5.4 Ablation Studies
Case study.
In addition, we conduct a case study on the hallucination case that may happen during the reasoning stage. For example, when being instructed to complete the incomplete triple (<missing-entity>, /olympics/olympic_games/parti-cipating_countries, /m/04vjh [Mauritania]) which queries the Olympic Games in which Mauritania has participated, the pre-trained LLM (GPT) incorrectly responds “Mauritania has never participated in the Olympic Games”. This error highlights how hallucinations can arise from the model’s lack of critical factual knowledge.
Nevertheless, triple (/m/04vjh [Mauritania], /olympics/olympic_participating_country/athletes - /olympics/olympic_athlete_affiliation/Olympics, /m/06sks6 [2012 Summer Olympics]) shows that Mauritania does attend 2012 Summer Olympic games, which can help the LLM to figure out Mauritania also participated in [2008 Summer Olympics] during the re-ranking stage. This case study also shows that access to accurate factual knowledge is essential for LLMs to perform reasoning tasks like KGC.
@@
Settings MRR Hits@1 Hits@3 Hits@10
Reordering (RotatE) 0.382 0.293 0.417 0.559
$\text{KGR}^{\mathbf{3}}$ (RotatE) 0.456 0.400 0.476 0.569
Reordering (GIE) 0.391 0.301 0.426 0.573
$\text{KGR}^{\mathbf{3}}$ (GIE) 0.463 0.400 0.485 0.581
Table 5: Ablation Experiments on FB15k-237 dataset with different re-ranking strategies.
Figure 3: Hits@$1$ performance over the amount of SFT data on the two datasets with different LLMs.
RQ5: How does the ranking strategy affect the performance of $\text{KGR}^{\mathbf{3}}$?
Rather than instructing the LLM to “sort” or “reorder” the entire candidate entity list, $\text{KGR}^{\mathbf{3}}$ guides the LLM to identify and place the most likely candidate at the top, while preserving the relative order of the remaining candidates. As shown in Table 5, we conduct additional experiments to compare these two ranking strategies. Under the reorder setting, the LLM is fine-tuned to follow the original ranking produced by the corresponding base KGC models. However, the results remain suboptimal, suggesting that LLMs are not well-suited for sorting entire entity lists. This limitation can be attributed to the nature of KGs, which do not inherently provide ground truth rankings of entities for each query triple. As a result, it is not intuitive to claim that ground truth tail entity A is more relevant than ground truth tail entity B for a given query $(h,r,?)$. Additionally, the experimental results demonstrate that constraining LLMs to mimic the ranking patterns of prior KGC methods does not alleviate their inherent limitations.
Furthermore, in some cases, LLMs may select an incorrect, or a false negative candidate answer, which reduces the rankings of the ground truth entity. For example, during the re-ranking stage, the Llama3-8B model suggests “Solihull” as the answer of query triple (England, location contains, <missing-entity>), where the ground truth is “Pontefract”. In fact, both “Solihull” and “Pontefract” are towns in England. Due to the inherent defect of the KG, the former is judged as a wrong answer. The re-ranking strategy adopted in $\text{KGR}^{3}$ ensures that when an LLM makes an incorrect judgment, the ranking of the ground truth answer drops by no more than 1 position, and hence, results in desirable performance improvements.
RQ6: How does the volume of SFT data affect the performance of $\text{KGR}^{\mathbf{3}}$?
@@ 6 Conclusion
Acknowledgement
The work described in this paper was partially supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 14222922, RGC GRF 2151185). We would also like to express our sincere gratitude to the reviewers and area chairs for their constructive comments and suggestions.